Homework: Homework #2 Score: 0 of 1 pt 4 of 15 (3 Problem 4.13 can see Heart Transplants 45.0 50.0 55.0 56.0 58.0 a) Using exponential smoothing with a of 0.60 and the given forecast for year 1, the forecasts for years 2 through 6 are (round your responses to ane decimal place) 56.6 Forecast 41.0 43.4 474 2 ound yo response to one decimal place For the forecast made us ng exponential smoothing with α-Ο 60 and the given forecast for year 1 MAD-52 surgeries Using ex ential smoothing with α of 0.90 and the given forecast for year 1, the forecasts for years 2 trough 6 are round your responses to one decinalplace Forecast 41.0 44.6 49.5 54.5 55.9 57.8 en forecast for year 1 MAD = 3.7 surgeries round your response to one decimal place For the forecast made using exponential smoothing with α·0.90 and the b) Forecasts for years 4 through 6 using a 3-year moving average are (round your responses to one decimal place) Forecast 5053.7 56.3 to one
Forecas 41.0 446 40.5 545 56.9 578 For the forecast made using expone tial smoothing with α·0 90 and the given forecast for year 1 MAD- surgeries ound your response to one decimal pla b) Forecasts for years 4 through 6 using a 3-year moving average are (round your responses to one decimal place): Forecast 50 53.7 56.3 to one Forecast 2 parts Clear All MacBook Pro 名0 F3 DW F4 2 4
Answer #1

Using regression

SUMMARY OUTPUT
Regression Statistics
Multiple R 0.9613
R Square 0.9242
Adjusted R Square 0.8989
Standard Error 1.6733
Observations 5
ANOVA
df SS MS F Significance F
Regression 1 102.4 102.4 36.5714 0.0091
Residual 3 8.4 2.8
Total 4 110.8
Coefficients Standard Error t Stat P-value Lower 95%
Intercept 43.2000 1.7550 24.6155 0.0001 37.6148
t 3.2000 0.5292 6.0474 0.0091 1.5160

y^ = 43.2 + 3.2 x

Observation Predicted y
1 46.4
2 49.6
3 52.8
4 56
5 59.2

for 6

y^ = 43.2 + 3.2*6 = 62.4